OpenStudy (anonymous):

Let n(A) = m and n(B)=n . if A has 15360 more subset than B, then (5m+3n) : (5m-3n) is as?

OpenStudy (anonymous):

@dpaInc @FoolForMath @apoorvk @eigenschmeigen @experimentX @mathmagician

OpenStudy (experimentx):

Looks like i forgot the basics .. is {1, 2} same as {2, 1}??

OpenStudy (anonymous):

@experimentX yea think so

OpenStudy (anonymous):

subsets*

OpenStudy (anonymous):

n(A) = m , n(B) = n number of subsets of set S where n(S) = s is 2^s i think

OpenStudy (anonymous):

yes the NO of substet = 2^n @eigenschmeigen u r correct go on

OpenStudy (experimentx):

oh .. kinds looks like \[ \sum_{i=1}^{n} \binom{n}{i} - \sum_{i=1}^{m} \binom{m}{i} = 15360\]

OpenStudy (anonymous):

2^m = 15360 + 2^n

OpenStudy (anonymous):

n, m are integers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

m = 14 , n = 10 works

OpenStudy (anonymous):

will there be other solutions?

OpenStudy (anonymous):

@eigenschmeigen we need to find ratio

OpenStudy (anonymous):

yes but if we solve then the ratio is trivial :D

OpenStudy (anonymous):

5:2

OpenStudy (anonymous):

but how u solved this plz show 2^m = 15360 + 2^n

OpenStudy (anonymous):

are there any other solutions for n and m though is what i'm interested in

OpenStudy (anonymous):

i kinda just saw that 15360 + 1024 = 16384

OpenStudy (anonymous):

i'm trying to algebraically solve it right now

OpenStudy (anonymous):

no u r correct....

OpenStudy (anonymous):

but there may be other solutions as far as we know

OpenStudy (anonymous):

no matter but how u solved this plz show 2^m = 15360 + 2^n

OpenStudy (anonymous):

any way plz show

OpenStudy (anonymous):

that's what i'm saying, i didnt really i just looked at it and figured it out. now i have a solution though

OpenStudy (anonymous):

ok then show it let us check that!

OpenStudy (anonymous):

notice that prime factorisation of 15360: \[15360 = 2^{10} \times 3 \times 5\] \[2^m = 15360 + 2^n\]\[2^m - 2^n = 15360 \]\[2^n(2^{m-n} -1) = 15360 \]\[2^n(2^{m-n} -1) = 2^{10} \times 3 \times 5 \] \[2^n(2^{m-n} -1) = 15360 = 2^{10} \times 15\]\[2^n(2^{m-n} -1) = 2^{10} (16 -1)\]\[2^n(2^{m-n} -1) = 2^{10} (2^{4} -1)\] so one solution is given by n = 10 , m-n = 4 m = 14

OpenStudy (anonymous):

i'm not sure whether i have proved this is unique, hopefully someone can confirm

OpenStudy (anonymous):

yes u r correct this was the way my teacher i just forfot to take a note at that time thxxx

OpenStudy (anonymous):

to show uniqueness it is necessary for \[2^k \times 15 \neq (2^q - 1) \text{ , } 0<k \le 10\]

OpenStudy (anonymous):

which is true as \[2^k(2^4 -1) = 2^{k+4} - 2^k\]